OpenStudy (anonymous):

How do solve this? X sqrt 64 x^3 + 2x sqrt x+8 sqrt 2x

OpenStudy (anonymous):

\[ x \sqrt{64x3} + 2x \sqrt{x} + 8 \sqrt{2x}\]